OpenStudy (anonymous):

x^3-x^2-16x+16 c=-4 is a zero What are the remaining zeros? What is this in full factored form?

OpenStudy (anonymous):

factor as \[x^3-x^2-16x+16 =(x+4)(\text{something})\] then use the quadratic formula. you can find the other factor by dividing, by synthetic division or by thinking.

OpenStudy (nenadmatematika):

x^2(x-1)-16(x-1)=(x-1)(x^2-16)=(x-1)(x-4)(x+4)

OpenStudy (nenadmatematika):

so the remaining zeros are 1, 4 and -4